OpenStudy (anonymous):

Δ ABC is isosceles. Angles B and C are congruent. The m∠A = 40°, m∠B = (3x + 1)°. Find x.

OpenStudy (anonymous):

you know, when it a isosceles triangle, which makes 2 angles equal to the third, right?

OpenStudy (anonymous):

Yeah?

OpenStudy (anonymous):

so you got this right? m<A=m<B right? so from here you can solve it :) it just gonna give you super ugly number though :)

OpenStudy (anonymous):

Are you sure? The answers are A) 13 B) 23 C) 46.3 D) 70 They aren't ugly?

OpenStudy (anonymous):

well,

OpenStudy (anonymous):

if there were your answers then you should do this: m<A + m<B = 180 ( we don't need C since B = C) 40 + 3x +1 =180 41 +3x =180 3x= 139 x =46.3 :)

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

cuz if you use m<A = m<B is gonna give you 13. something ( very ugly)

OpenStudy (anonymous):